# The Atlas Model and SDEs with Boundary Interaction

### v1

@@ 2.1 Tightness of the Atlas Model
Let $B^{N}=(B^{N}_{t})_{t\geq 0}$ be the unique solution to SDE 1.2 for $N\geq 1$ and let $(X^{i}-B^{N},L^{i})$ be the solution to the Skorokhod problem for $\xi_{i}+W^{i}-B^{N}$ for $i\in[N]$. Whenever we wish to emphasise the size of the particle system, we will add the number $N$ to the superscript of $X^{i}$, $L^{i}$, and other particle-related quantities. Recall that we defined the random measures $\beta^{N}$ on $0,\infty)\times\mathbb{R}$ by $\mathrm{d}\beta^{N}(t,x)=\mathrm{d}\delta_{B^{N}_{t}}\mathrm{d}t$. Our first goal is to prove that for any $i\in\mathbb{N}$, the sequence $(L^{N,i},\beta^{N})_{N\geq i}$ is tight on $C([0,\infty))\times\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$. Recall that the space $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$ of measures $m$ on $0,\infty)\times\mathbb{R}$ such that $m([0,t]\times\mathbb{R})=t$ for all $t\geq 0$ is endowed with the initial topology generated by the projections $\mathcal{M}_{1}([0,\infty)\to\mathcal{P}([0,n]\times\mathbb{R})$, $m\mapsto\frac{1}{n}m|_{[0,n]\times\mathbb{R}}$ for $n\in\mathbb{N}$. Note that by the properties of the initial topology, in order to show that a subset of $\mathcal{M}_{1}([0,\infty)$ is compact, it suffices to show that the image of the subset under all the projections is compact. From this it follows that a sequence of random variables with values in $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$ is tight if each projection of the sequence is tight.
@@
Tightness of $(L^{N,i})_{N\geq i}$: It suffices to show that for any $T>0$, the restriction of $L^{N,i}$, $N\geq i$, to the interval $[0,T]$ is tight on $C([0,T])$. So let us fix $T>0$ and denote by $\mathcal{T}^{N}_{[0,T]}$ the set of $\mathbb{F}^{N}$-stopping times with values in $[0,T]$. Appealing to Aldous’ tightness criterion (see e.g. [6, Theorem 16.10]), we must show that for all $\eta>0$,
@@
The idea of the proof is as follows: first, we prove that for any $N\geq 1$ and any time $\tau\in\mathcal{T}^{N}_{[0,T]}$, there is a positive fraction of particles near $X^{N,i}_{\tau}$. Then, if $L^{N,i}$ were to grow rapidly between $\tau$ and $\tau+\theta$, all particles $j\in[N]$ sufficiently close to $X^{N,i}_{\tau}$ at time $\tau$ would be pushed upwards by an increase in their corresponding regulator process $L^{N,j}$. This in turn would contradict the fact that the empirical average of the regulators growth at a finite rate $\gamma$. Thus, the increments of $L^{N,i}$ cannot be too large as quantified by (2.3).
@@
Fix $j\in[N]$ and set $Z^{N,j}_{t}=\lvert X^{N,j}_{t}-X^{N,i}_{t}\rvert$. By the Itô–Tanaka formula (see [23, Chapter 3, Theorem 7.1]), we have that
@@
with reflection at the origin. The process $L^{N,j}+L^{N,i}$ is viewed as an exogenous input to this SDE. By standard comparison results for reflected SDE, the process $Z^{N,j}$ is dominated by the solution $\tilde{Z}^{N,j}$ to the SDE
@@
Tightness of $(\beta^{N})_{N\geq 1}$: Next, let us consider the sequence $(\beta^{N})_{N\geq 1}$. By the remark immediately before the statement of the proposition, we have to show that for any $n\in\mathbb{N}$, the family $\frac{1}{n}\beta^{N}|_{[0,n]\times\mathbb{R}}$, $n\geq 1$, is tight on $\mathcal{P}([0,n]\times\mathbb{R})$. Owing to [35, Proposition 2.2(ii)], this is true if the family of measures $\mathcal{B}([0,n]\times\mathbb{R})\ni A\mapsto\frac{1}{n}\mathbb{E}[\beta^{N}(A)]$, $N\geq 1$ is tight. This in turn follows if we can show that
$\mathbb{E}\biggl[\int_{[0,n]\times\mathbb{R}}\lvert x\rvert\,\mathrm{d}\beta^{N}(t,x)\biggr]=\mathbb{E}\biggl[\int_{0}^{n}\lvert B^{N}_{t}\rvert\,\mathrm{d}t\biggr]$
is bounded uniformly over $N\geq 1$. Let us note that
$\int_{0}^{n}\lvert B^{N}_{t}\rvert\,\mathrm{d}t=\int_{0}^{n}(B^{N}_{t})_{-}+(B^{N}_{t})_{+}\,\mathrm{d}t\leq\int_{0}^{n}(B^{N}_{t})_{-}\,\mathrm{d}t+\int_{0}^{n}(X^{N,1}_{t})_{+}\,\mathrm{d}t.$ (2.14)
Since
$\mathbb{E}[\lvert X^{N,1}_{t}\rvert]\leq\mathbb{E}[\xi_{1}]+\mathbb{E}[\lvert W^{1}_{t}\rvert]+\mathbb{E}\biggl[\int_{0}^{t}\gamma N\mathbf{1}_{\{X^{N,1}_{s}=B^{N}_{s}\}}\,\mathrm{d}s\biggr]\leq\mathbb{E}[\xi_{1}]+\frac{\sqrt{2t}}{\sqrt{\pi}}+\gamma t,$
the second term on the right-hand side of (2.14) is clearly bounded uniformly in $N$. Thus, it remains to derive a suitable bound for the first term. To achieve this, we introduce an auxiliary system whose initial conditions are given by $\tilde{\xi}_{j}=\xi_{j}\land 1$, $j\in[N]$. Let $\tilde{X}^{1}$, …, $\tilde{X}^{N}$ denote the corresponding solution of the Atlas model and define $\tilde{B}^{N}=(\tilde{B}^{N}_{t})_{t\geq 0}$ by $\tilde{B}^{N}_{t}=\min_{j\in[N]}\tilde{X}^{j}_{t}$, so that $\tilde{B}^{N}$ solves the reflected SDE (1.2). By Proposition 2.2, we have that $\tilde{B}^{N}_{t}\leq B^{N}_{t}$, so that $(B^{N}_{t})_{-}\leq(\tilde{B}^{N}_{t})_{-}$. Now, Itô’s formula yields
@@
We take expectations on both sides of the above and use that $\tilde{B}^{N}_{t}\leq-(\tilde{B}^{N}_{t})_{-}+(\tilde{X}^{N,1}_{t})_{+}\leq-(\tilde{B}^{N}_{t})_{-}+\lvert X^{N,1}_{t}\rvert$ and $\frac{1}{N}\sum_{j=1}^{N}\lvert\tilde{X}^{j}_{n}\rvert^{2}\geq 0$ to arrive at
$2\gamma\mathbb{E}\biggl[\int_{0}^{n}(B^{N}_{t})_{-}\,\mathrm{d}t\biggr]\leq 2\gamma\mathbb{E}\biggl[\int_{0}^{n}(\tilde{B}^{N}_{t})_{-}\,\mathrm{d}t\biggr]\leq 1+n+\mathbb{E}\biggl[\int_{0}^{n}\lvert X^{N,1}_{t}\rvert\,\mathrm{d}t\biggl].$
We already saw that the right-hand side above is bounded uniformly over $N\geq 1$, so we obtain the desired uniform bound over $\mathbb{E}[\int_{0}^{n}\lvert B^{N}_{t}\rvert\,\mathrm{d}t]$.
Moment bound for $(L^{N,i})_{N\geq i}$: To prove the final statement (2.2), for $j\in[N]$ and $t\geq 0$, we write
$\lvert L^{N,j}_{t}\rvert^{2}$ $=2\int_{0}^{t}L^{N,j}_{s}\,\mathrm{d}L^{N,j}_{s}$
$=2\int_{0}^{t}X^{N,j}_{s}\,\mathrm{d}L^{N,j}_{s}-2\int_{0}^{t}(\xi_{j}+W^{j}_{s})\,\mathrm{d}L^{N,j}_{s}$
$\leq 2\int_{0}^{t}B^{N}_{s}\,\mathrm{d}L^{N,j}_{s}-2W^{j}_{t}L^{N,j}_{t}+2\int_{0}^{t}L^{N,j}_{s}\,\mathrm{d}W^{j}_{t}$
$\leq 2\int_{0}^{t}B^{N}_{s}\,\mathrm{d}L^{N,j}_{s}+2\lvert W^{j}_{t}\rvert^{2}+\frac{1}{2}\lvert L^{N,j}_{t}\rvert^{2}+2\int_{0}^{t}L^{N,j}_{s}\,\mathrm{d}W^{i}_{t}.$
@@
We rearrange this inequality, sum over $j\in[N]$, and take expectation on both sides to get
@@
$\mathbb{E}\bigl[\lvert L^{N,i}_{t}\rvert^{2}\bigr]=\frac{1}{N}\sum_{j=1}^{N}\mathbb{E}\bigl[\lvert L^{N,j}_{t}\rvert^{2}\bigr]\leq 4\gamma\mathbb{E}\biggl[\int_{0}^{t}B^{N}_{s}\,\mathrm{d}s\biggr]+4t.$
From our previous observations, we know that $\mathbb{E}[\int_{0}^{t}B^{N}_{s}\,\mathrm{d}s]$ is bounded uniformly over $N\geq i$, so we can conclude the same for $\mathbb{E}[\lvert L^{N,i}_{t}\rvert^{2}]$. ∎

### v2

@@ 2.1 Tightness of the Atlas Model
Let $B^{N}=(B^{N}_{t})_{t\geq 0}$ be the unique solution to SDE (1.2) for $N\geq 1$ and let $(X^{i}-B^{N},L^{i})$ be the solution to the Skorokhod problem for $\xi_{i}+W^{i}-B^{N}$ for $i\in[N]$. Whenever we wish to emphasise the size of the particle system, we will add the number $N$ to the superscript of $X^{i}$, $L^{i}$, and other particle-related quantities. Recall that we defined the random measures $\beta^{N}$ on $0,\infty)\times\mathbb{R}$ by
$\mathrm{d}\beta^{N}(t,x)=\mathrm{d}\delta_{B^{N}_{t}}\mathrm{d}t.$
Our goal is to prove that for any $i\in\mathbb{N}$, the sequence $(L^{N,i},\beta^{N})_{N\geq i}$ is tight on $C([0,\infty))\times\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$. Here the space $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$ of measures $m$ on $0,\infty)\times\mathbb{R}$ such that $m([0,t]\times\mathbb{R})=t$ for all $t\geq 0$ is endowed with the initial topology generated by the projections $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})\to\mathcal{P}([0,n]\times\mathbb{R})$, $m\mapsto\frac{1}{n}m|_{[0,n]\times\mathbb{R}}$ for $n\in\mathbb{N}$. Note that by the properties of the initial topology, in order to show that a subset of $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$ is compact, it suffices to show that the image of the subset under all the projections is compact. From this it follows that a sequence of random variables with values in $\mathcal{M}_{1}([0,\infty)\times\mathbb{R})$ is tight if each projection of the sequence is tight.
@@
Tightness of $(L^{N,i})_{N\geq i}$: It suffices to show that for any $T>0$, the restriction of $L^{N,i}$, $N\geq i$, to the interval $[0,T]$ is tight on $C([0,T])$. So let us fix $T>0$ and denote by $\mathcal{T}^{N}_{[0,T]}$ the set of $\mathbb{F}^{N}$-stopping times with values in $[0,T]$. Appealing to Aldous’ tightness criterion (see e.g. [undefe, Theorem 16.10]), we must show that for all $\eta>0$,
@@
The idea of the proof is as follows: first, we prove that for any $N\geq 1$ and any time $\tau\in\mathcal{T}^{N}_{[0,T]}$, there is a positive fraction of particles near $X^{N,i}_{\tau}$. Then, if $L^{N,i}$ were to grow rapidly between $\tau$ and $\tau+\theta$, all particles $j\in[N]$ sufficiently close to $X^{N,i}_{\tau}$ at time $\tau$ would be pushed upwards by an increase in their corresponding regulator process $L^{N,j}$. This in turn would contradict the fact that the empirical average of the regulators growth at the finite rate $\gamma$. Thus, the increments of $L^{N,i}$ cannot be too large as quantified by (2.4).
@@
Fix $j\in[N]$ and set $Z^{N,j}_{t}=\lvert X^{N,j}_{t}-X^{N,i}_{t}\rvert$. By the Itô–Tanaka formula (see [undefv, Chapter 3, Theorem 7.1]), we have that
@@
with reflection at the origin. The process $L^{N,j}+L^{N,i}$ is viewed as an exogenous input to this SDE. By standard comparison results for reflected SDEs, the process $Z^{N,j}$ is dominated by the solution $\tilde{Z}^{N,j}$ to the SDE
@@
Tightness of $(\beta^{N})_{N\geq 1}$: Next, let us consider the sequence $(\beta^{N})_{N\geq 1}$. By the remark immediately before the statement of the proposition, we have to show that for any $n\in\mathbb{N}$, the family $\frac{1}{n}\beta^{N}|_{[0,n]\times\mathbb{R}}$, $n\geq 1$, is tight on $\mathcal{P}([0,n]\times\mathbb{R})$. Owing to [undefah, Proposition 2.2(ii)], this is true if the family of probability measures $\mathcal{B}([0,n]\times\mathbb{R})\ni A\mapsto\frac{1}{n}\mathbb{E}[\beta^{N}(A)]$, $N\geq 1$, is tight. This in turn follows if we can show that for every $\epsilon>0$ there exists $K>0$ such that
$\mathbb{E}\Bigl[\beta^{N}\Bigl([0,n]\times[-K,K]^{c}\Bigr)\Bigr]=\mathbb{E}\biggl[\int_{0}^{n}\mathbf{1}_{\{\lvert B^{N}_{t}\rvert>K\}}\,\mathrm{d}t\biggr]=\int_{0}^{n}\mathbb{P}(\lvert B^{N}_{t}\rvert>K)\,\mathrm{d}t<\epsilon$
for all $N\geq 1$. Let us note that
$\lvert B^{N}_{t}\rvert=(B^{N}_{t})_{-}+(B^{N}_{t})_{+}\leq(B^{N}_{t})_{-}+(X^{N,1}_{t})_{+}\leq(B^{N}_{t})_{-}+\lvert X^{N,1}_{t}\rvert$,
so that
$\int_{0}^{n}\mathbb{P}(\lvert B^{N}_{t}\rvert>K)\,\mathrm{d}t\leq\int_{0}^{n}\mathbb{P}\bigl((B^{N}_{t})_{-}>K/2\bigr)\,\mathrm{d}t+\int_{0}^{n}\mathbb{P}\bigl(\lvert X^{N,1}_{t}\rvert>K/2\bigr)\,\mathrm{d}t.$ (2.15)
We shall bound the two terms on the right-hand side above separately. For the second one, we have
$\mathbb{P}\bigl(\lvert X^{N,1}_{t}\rvert>K/2\bigr)$ $\leq\mathbb{P}\bigl(\lvert\xi_{1}\rvert>K/6\bigr)+\mathbb{P}\bigl(\lvert W^{1}_{t}\rvert>K/6\bigr)+\mathbb{P}\bigl(\lvert L^{N,1}_{t}\rvert>K/6\bigr)$
$\leq\mathbb{P}\bigl(\lvert\xi_{1}\rvert>K/6\bigr)+\mathbb{P}\bigl(\lvert W^{1}_{n}\rvert>K/6\bigr)+\frac{6\mathbb{E}[L^{N,1}_{n}]}{K}$
$\leq\mathbb{P}\bigl(\lvert\xi_{1}\rvert>K/6\bigr)+\mathbb{P}\bigl(\lvert W^{1}_{n}\rvert>K/6\bigr)+\frac{6\gamma n}{K}$
Clearly, by choosing $K>0$ large enough, we can make the right-hand side above smaller than $\frac{\epsilon}{2n}$, so that
$\int_{0}^{n}\mathbb{P}\bigl(\lvert X^{N,1}_{t}\rvert>K/2\bigr)\,\mathrm{d}t<\frac{\epsilon}{2}.$ (2.16)
To bound the first integral on the right-hand side of (2.15), we shall appeal to Proposition 2.2 (i). We introduce an auxiliary system whose initial conditions are given by $\tilde{\xi}_{j}=\xi_{j}\land 1$, $j\in[N]$. Let $\tilde{X}^{1}$, …, $\tilde{X}^{N}$ denote the corresponding solution of the Atlas model and define $\tilde{B}^{N}=(\tilde{B}^{N}_{t})_{t\geq 0}$ by $\tilde{B}^{N}_{t}=\min_{j\in[N]}\tilde{X}^{j}_{t}$, so that $\tilde{B}^{N}$ solves the reflected SDE (1.2). By Proposition 2.2 (i), we have that $\tilde{B}^{N}_{t}\leq B^{N}_{t}$, so that $(B^{N}_{t})_{-}\leq(\tilde{B}^{N}_{t})_{-}$. Now, Itô’s formula yields
@@
We take expectations on both sides of the above and use that $(\tilde{B}^{N}_{t})_{-}=-\tilde{B}^{N}_{t}+(\tilde{B}^{N}_{t})_{+}\leq-\tilde{B}^{N}_{t}+(\tilde{X}^{1}_{t})_{+}$ and $\frac{1}{N}\sum_{j=1}^{N}\lvert\tilde{X}^{j}_{n}\rvert^{2}\geq 0$ to arrive at
$2\gamma\mathbb{E}\biggl[\int_{0}^{n}(B^{N}_{t})_{-}\,\mathrm{d}t\biggr]$ $\leq 2\gamma\mathbb{E}\biggl[\int_{0}^{n}(\tilde{B}^{N}_{t})_{-}\,\mathrm{d}t\biggr]$
$\leq 1+n+\mathbb{E}\biggl[\int_{0}^{n}(\tilde{X}^{1}_{t})_{+}\,\mathrm{d}t\biggl]$
$\leq 1+2n+\frac{2\sqrt{2}}{3\sqrt{\pi}}n^{3/2}+\gamma\frac{n^{2}}{2}.$ (2.17)
But since
$\int_{0}^{n}\mathbb{P}\bigl((B^{N}_{t})_{-}>K/2\bigr)\,\mathrm{d}t\leq\frac{2}{K}\mathbb{E}\biggl[\int_{0}^{n}(B^{N}_{t})_{-}\,\mathrm{d}t\biggr],$
the above implies that by enlarging $K$ if necessary, we have $\int_{0}^{n}\mathbb{P}\bigl((B^{N}_{t})_{-}>K/2\bigr)\,\mathrm{d}t<\frac{\epsilon}{2}$. Together with (2.16), we get $\int_{0}^{n}\mathbb{P}(\lvert B^{N}_{t}\rvert>K)\,\mathrm{d}t<\epsilon$ as required.
Moment bound for $(L^{N,i})_{N\geq i}$: To prove the final statement (2.3), we combine Proposition 2.2 (ii) and (iii). Define the indicators $I_{j}=\mathbf{1}_{\{\xi_{j}\leq 1\}}$ for $j\in[N]$. By Proposition 2.2 (ii), it holds that
$\mathbb{E}\bigl[\lvert L^{N,j}_{t}\rvert^{2}\big|I_{j}=1\bigr]\geq\mathbb{E}\bigl[\lvert L^{N,j}_{t}\rvert^{2}\big|I_{j}=0\bigr],$
which in view of the law of total probability implies that
$\mathbb{E}\bigl[\lvert L^{N,j}_{t}\rvert^{2}\bigr]\leq\mathbb{E}\bigl[\lvert L^{N,j}_{t}\rvert^{2}\big|I_{j}=1\bigr].$ (2.18)
Next, consider the system consisting of the particles in $\mathcal{I}_{N}=\{j\in[N]\mathpunct{:}I_{j}=1\}$. On $\{\mathcal{I}_{N}\neq\varnothing\}$, we can solve SDE (1.2) with initial conditions $\xi_{j}$, $j\in\mathcal{I}_{N}$. Denote the solution by $\tilde{B}^{N}$ and let $(\tilde{X}^{N,j}-\tilde{B}^{N},\tilde{L}^{N,j})$, $j\in\mathcal{I}_{N}$, be the solution to the Skorokhod problem for $\xi_{j}+W^{j}-\tilde{B}^{N}$. Note that these processes should be distinguished from the ones we constructed when proving tightness of $(\beta^{N})_{N\geq 1}$.
Owing to Proposition 2.2 (iii), on $\{\mathcal{I}_{N}\neq\varnothing\}$, we have $\tilde{L}^{N,j}_{t}\geq L^{N,j}_{t}$ for $j\in\mathcal{I}_{N}$. From this and (2.18), we deduce
$\mathbb{E}\bigl[\lvert L^{N,i}_{t}\rvert^{2}\bigr]\leq\mathbb{E}\bigl[\lvert L^{N,i}_{t}\rvert^{2}\big|I_{i}=1\bigr]\leq\mathbb{E}\bigl[\lvert\tilde{L}^{N,i}_{t}\rvert^{2}\big\rvert I_{i}=1\bigr]=\frac{\mathbb{E}\bigl[I_{i}\lvert\tilde{L}^{N,i}_{t}\rvert^{2}\bigr]}{\mathbb{E}[I_{i}]}.$ (2.19)
Thus, we are done if we can show that $\mathbb{E}[I_{i}\lvert\tilde{L}^{N,i}_{t}\rvert^{2}]$ is bounded uniformly over $N\geq i$. For $j\in\mathcal{I}_{N}$ and $t\geq 0$, let us write
$\lvert\tilde{L}^{N,j}_{t}\rvert^{2}$ $=2\int_{0}^{t}\tilde{L}^{N,j}_{s}\,\mathrm{d}\tilde{L}^{N,j}_{s}$
$=2\int_{0}^{t}\tilde{X}^{N,j}_{s}\,\mathrm{d}\tilde{L}^{N,j}_{s}-2\int_{0}^{t}(\xi_{j}+W^{j}_{s})\,\mathrm{d}\tilde{L}^{N,j}_{s}$
$\leq 2\int_{0}^{t}\tilde{B}^{N}_{s}\,\mathrm{d}\tilde{L}^{N,j}_{s}-2W^{j}_{t}\tilde{L}^{N,j}_{t}+2\int_{0}^{t}\tilde{L}^{N,j}_{s}\,\mathrm{d}W^{j}_{s}$
$\leq 2\int_{0}^{t}\tilde{B}^{N}_{s}\,\mathrm{d}\tilde{L}^{N,j}_{s}+2\lvert W^{j}_{t}\rvert^{2}+\frac{1}{2}\lvert\tilde{L}^{N,j}_{t}\rvert^{2}+2\int_{0}^{t}\tilde{L}^{N,j}_{s}\,\mathrm{d}W^{i}_{s}.$
@@
We rearrange this inequality, multiply both sides by $I_{j}$, sum over $j\in\mathcal{I}_{N}$, and take expectation to get
@@
$\mathbb{E}\bigl[I_{i}\lvert\tilde{L}^{N,i}_{t}\rvert^{2}\bigr]=\frac{1}{N}\sum_{j=1}^{N}\mathbb{E}\bigl[I_{j}\lvert\tilde{L}^{N,j}_{t}\rvert^{2}\bigr]\leq 4\gamma\mathbb{E}\biggl[\mathbf{1}_{\{\mathcal{I}_{N}\neq\varnothing\}}\int_{0}^{t}\tilde{B}^{N}_{s}\,\mathrm{d}s\biggr]+4t,$
where we used that
$\sum_{j=1}^{N}I_{j}\tilde{L}^{N,j}_{s}=\sum_{j\in\mathcal{I}_{N}}\tilde{L}^{N,j}_{s}=\gamma s$
for $s\geq 0$ on $\{\mathcal{I}_{N}\neq\varnothing\}$. Now, analogously to the previous step, it can be shown that the expression $\mathbb{E}[\mathbf{1}_{\{\mathcal{I}_{N}\neq\varnothing\}}\int_{0}^{t}\tilde{B}^{N}_{s}\,\mathrm{d}s]$ is bounded uniformly over $N\geq i$ and so the same is true for $\mathbb{E}[I_{i}\lvert\tilde{L}^{N,i}_{t}\rvert^{2}]$. Together with (2.19) this concludes the proof. ∎